OpenStudy (anonymous):

Help please!!! Which of the following is a solution of x2 - 7x = -5? 7 plus or minus the square root of 69 divided by two 7 plus or minus the square root of 29 divided by two. negative 7 plus or minus the square root of 29 divided by two negative 7 plus or minus the square root of 69 divided by two.

hartnn (hartnn):

do you know quadratic formula ?

OpenStudy (anonymous):

Yes, b+/-sqrt(b^2-2ac)/2a

hartnn (hartnn):

so, we have x^2 - 7x +5 = 0 can you find a,b,c ?

OpenStudy (anonymous):

It's C, right?

hartnn (hartnn):

i didn't work out yet, did you work it out ?

OpenStudy (anonymous):

because, it would be -7+/-sqrt(-7^2-2(1)(-5))/2(1)

hartnn (hartnn):

first of all, b = -7 so, -b = +7

hartnn (hartnn):

so its not C or D

OpenStudy (anonymous):

Riiighhht, I forgot about that part, damn. I think I have the formula wrong then.

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

Okay, i've got it! It's 7+/-sqrt(29)/2 so the answer is B!

hartnn (hartnn):

b = -7 a = 1 c = 5 b^2 - 4ac = 49 - 20 = 29 yes, you're correct :)

OpenStudy (anonymous):

Can you help me with another one if possible(?)

hartnn (hartnn):

sure

OpenStudy (anonymous):

I already know it's not A. But, I don't know which one of the other 3 it could be. Solve x2 + 6x + 7 = 0. x = -1 and x = -5 3 plus or minus square root of 2 negative 3 plus or minus square root of 2 quantity of negative 3 plus or minus square root of 2 all over 2

hartnn (hartnn):

what values of a,b,c you got ?

OpenStudy (anonymous):

1, 6 and 7.

hartnn (hartnn):

yes b^2 -4ac =....

OpenStudy (anonymous):

6^2-4(1)(7) 36-28 8

hartnn (hartnn):

good, so where are you stuck ?? proceed

hartnn (hartnn):

[-6 +/- sqrt 8 ] /2 got this ?

OpenStudy (anonymous):

Yeah. But then what, how does it end up like that?

hartnn (hartnn):

-6/2 +/- (sqrt 8)/2 so you need to simplify sqrt 8 can you factor 8 ?

OpenStudy (anonymous):

There's 4 and 2.

hartnn (hartnn):

good! \(\sqrt 8 = \sqrt 4 \sqrt 2\) whats sqrt 4 ??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so it would go as 2sqrt(2)

hartnn (hartnn):

yes

hartnn (hartnn):

-6/2 +/- (sqrt 8)/2 =-6/2 +/- ( 2 sqrt 2)/2 simplify ?

OpenStudy (anonymous):

(-3+/-(2sqrt2))/2 so iiiittsss D!

hartnn (hartnn):

=-6/2 +/- ( 2 sqrt 2)/2 = -3 +/- sqrt 2 the '2' gets cancelled from numerator and denominator for BOTH the terms

hartnn (hartnn):

doubts ?

OpenStudy (anonymous):

So it just ends up just being C?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Thank you sooo much!!!

hartnn (hartnn):

welcome ^_^